# Jet production at electron-positron colliders at next-to-next-to-next-to-leading order in QCD

### v1

@@ Abstract
preprint: IPPP/25/27, ZU-TH 36/25
Introduction—Experiments at present and future colliders plan to deliver an unprecedented amount of data and significantly increase our sensitivity to new physics if equally precise Standard Model predictions are available. In high-energy collisions, the effects of QCD are dominant and their evaluation at higher perturbative orders in the strong coupling $\alpha_{s}$ requires techniques for the identification and removal of infrared (IR) divergences from the constituent scattering amplitudes, known as infrared subtraction schemes.
@@
General approaches have been developed for next-to-leading order (NLO) calculations [1, 2], followed by several strategies at next-to-next-to-leading order (NNLO) [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and ongoing efforts to include higher-multiplicity final states and to generalize the current formalisms [17, 18, 19, 20, 21]. NNLO-accurate predictions are nowadays available for several Standard Model processes, see [22] for a review. N3LO accuracy has recently been reached for color-singlet production and decay processes for both inclusive [23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and differential [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] observables, relying either on $q_{T}$-slicing [7] or the Projection-to-Born [11] method. Neither of these approaches is applicable to arbitrary processes and observables, especially in the presence of final-state jets, making the development of a general N3LO subtraction scheme highly desirable.
@@
This work represents the first N3LO differential calculation obtained with a subtraction scheme that is not restricted to specific classes of processes and does not rely on the prior knowledge of inclusive quantities. Our calculation is therefore a crucial step in tackling more complicated processes with jets at high perturbative order [57] relevant to the physics goals of the upcoming High-Luminosity phase of the LHC [58]. Moreover, results at this level of precision are relevant for future QCD precision studies [59] at the next generation of electron-positron colliders, such as the FCC-ee [60, 61, 62] and CEPC [63, 64, 65].
Method—In the following, we outline the extension of antenna subtraction for final-state radiation to N3LO. The $\mathcal{O}(\alpha_{s}^{3})$ correction to two-jet production at electron-positron colliders receives contributions from final states with up to five partons. We denote the contribution to the N3LO coefficient from a configuration with $n$ final-state partons by $\text{d}\sigma^{(3)}_{n}$. The triple-real contribution $\text{d}\sigma^{(3)}_{5}$ is given by three real emissions on top of the Born-level process, the double-real-virtual contribution $\text{d}\sigma^{(3)}_{4}$ by two real emissions at one loop, the real-double-virtual contribution $\text{d}\sigma^{(3)}_{3}$ by one real emission at two loops and finally the triple-virtual contribution $\text{d}\sigma^{(3)}_{2}$ by the three-loop correction. Examples of Feynman diagrams contributing to different types of N3LO corrections are shown in Figure 1.
The individual contributions contain explicit IR singularities due to loop corrections ($\text{d}\sigma^{(3)}_{2}$), implicit singularities arising when the external momenta approach soft and collinear limits ($\text{d}\sigma^{(3)}_{5}$), or a mixture of both ($\text{d}\sigma^{(3)}_{3}$ and $\text{d}\sigma^{(3)}_{4}$). These divergences cancel when virtual and real corrections are combined in physical predictions [66, 67], but each individual term is numerically integrable over the associated phase space only after introducing dedicated subtraction terms. We denote by $\text{d}\sigma^{(3)}_{n,\text{finite}}$ the individual contributions after subtraction, free from any type of IR singularity. The definition of $\text{d}\sigma^{(k)}_{n,\text{finite}}$ up to NNLO in the context of antenna subtraction was described in [3, 5]. At N3LO, we write schematically
$\text{d}\sigma_{5,\text{finite}}^{(3)}$ $=\text{d}\sigma_{5}^{(3)}$ $-\text{d}\sigma^{(3)}_{5,S_{1}}$ $-\text{d}\sigma^{(3)}_{5,S_{2}}$ $-\text{d}\sigma^{(3)}_{5,S_{3}}\,,$ (1)
$\text{d}\sigma_{4,\text{finite}}^{(3)}$ $=\text{d}\sigma_{4}^{(3)}$ $-\text{d}\sigma^{(3)}_{4,V_{1}}$ $-\text{d}\sigma^{(3)}_{4,V_{1}S_{1}}$ $-\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}\,,$ (2)
$\text{d}\sigma_{3,\text{finite}}^{(3)}$ $=\text{d}\sigma_{3}^{(3)}$ $-\text{d}\sigma^{(3)}_{3,V_{2}}$ $-\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}\,,$ (3)
@@
$\text{d}\sigma^{(3)}_{4,V_{1}}$ $=$ $-\int_{1}\text{d}\sigma^{(3)}_{5,S_{1}}\,,$ (5)
$\text{d}\sigma^{(3)}_{3,V_{2}}$ $=$ $-\int_{1}\text{d}\sigma^{(3)}_{4,V_{1}S_{1}}-\int_{2}\text{d}\sigma^{(3)}_{5,S_{2}}\,,$ (6)
$\text{d}\sigma^{(3)}_{2,V_{3}}$ $=$ $-\int_{1}\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}-\int_{2}\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}-\int_{3}\text{d}\sigma^{(3)}_{5,S_{3}}\,,$ (7)
@@
The subtraction terms are built from antenna functions with the following notation: $X_{n}^{\ell}$ is a $n$-parton $\ell$-loop antenna function intended to subtract the singular behavior of $n-2$ unresolved partons between a pair of hard radiators at $\ell$ loops. $X$ is a placeholder for a particular partonic configuration [3]. The quark-antiquark antenna functions relevant for our process are obtained from color-ordered matrix elements for the decay of a photon to $n$ partons [45, 70], denoted by $M_{n,\gamma^{*}}^{\ell}$. The integrated counterpart of an antenna function, $\mathcal{X}_{n}^{\ell}$, is obtained by analytic integration over the inclusive phase space of the $n-2$ unresolved emissions. All the N3LO final-state antenna functions were recently integrated in [70, 71, 72].
@@
where $M_{n}^{0}$ represents a generic tree-level color-ordered matrix element for $e^{+}e^{-}\to n$ partons and the shorthand symbol $\otimes$ encodes a momentum mapping from a higher-multiplicity to a lower-multiplicity phase space, in order to absorb unresolved emissions. For each term, a sum over external partons and color correlations is understood. The jet function $J_{n_{j}}^{(n_{p})}$ reconstructs $n_{j}$ jets from $n_{p}$ final-state partons. The sums in (13a) and (13b) run over configurations where the two unresolved emissions are respectively color-connected ($cc$) and almost color-connected ($acc$[3, 68].
@@
Here the $X_{5}^{0}$ antenna functions address color-connected triple-unresolved emissions. In particular, they contain the singular behavior in triple-soft [75, 76, 77] and quadruple-collinear [78, 79] limits. A $5\to 2$ antenna momentum mapping [80] is used to relate the triple-real and the fully resolved phase space. The remaining terms are needed to remove the single- and double-unresolved limits of the $X_{5}^{0}$, which are already subtracted by $\text{d}\sigma^{(3)}_{5,S_{1}}+\text{d}\sigma^{(3)}_{5,S_{2}}$. The five-particle antenna functions mimic the full matrix element in unresolved limits and factorize onto lower-multiplicity antenna functions as follows:
[…]
@@
Calculation—Our calculation was performed with a version of the NNLOJET Monte Carlo framework [81] extended for an N3LO calculation. For the N3LO antenna functions $X_{n}^{\ell}$, we adapted the implementation of the relevant matrix elements in EERAD3 [82]. The necessary helicity amplitudes for $e^{+}e^{-}\to\text{partons}$ are available in NNLOJET as part of the calculation of three-jet production at NNLO [83]. The five-parton tree-level amplitudes used in our calculation were computed in [84, 85, 86], the four-parton one-loop amplitudes in [87, 88, 89] and the three-parton two-loop amplitudes in [90, 91]. Finally, the three-loop quark form factor was computed in [92, 93, 94].
@@
Results—All our results are normalized by the respective Born-level cross section $\sigma_{0}$ for $e^{+}e^{-}\to q\bar{q}$. As a consequence, the electroweak coupling drops out and our results are valid for both $\gamma^{*}$ and $Z$ as the intermediate boson. We neglect the singlet contribution which is numerically small, as observed for example in [99, 100]. We set $\sqrt{s}=m_{Z}=91.2$ GeV and $\alpha_{s}(m_{Z})=0.118$. However, for the observables we consider, the dependence on the value of the energy scale only enters through the running of the strong coupling. The renormalization scale is chosen to be equal to the center-of-mass energy, $\mu_{r}=\sqrt{s}$, and theory uncertainties are estimated by varying it within $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$.
@@
$\sigma^{(3)}_{2j}(y_{\text{cut}}=0.1)$ $=\sigma_{0}\left(\frac{\alpha_{s}}{2\pi}\right)^{3}(-455\pm 32)\,,$ (19)
$\sigma^{(3)}_{2j}(y_{\text{cut}}=0.01)$ $=\sigma_{0}\left(\frac{\alpha_{s}}{2\pi}\right)^{3}(-2554\pm 57)\,.$ (20)
@@
Conclusions—We presented the first direct fully differential calculation of jet production in electron-positron annihilation at third order in QCD. We described the structure of the IR counterterms for the final-state radiation in $e^{+}e^{-}\to jj$ within antenna subtraction, including novel triple-unresolved limits. As an application, we computed the production rate of two jets and the distribution of the energy of the leading jet at N3LO at different values of the jet reconstruction parameter.
@@
Acknowledgements.
Acknowledgments—We are grateful to Thomas Gehrmann, Nigel Glover and Aude Gehrmann-de Ridder for support and encouragement to pursue this work. We thank Elliot Fox for comments on the manuscript. XC is supported by the National Science Foundation of China (NSFC) with grant No.12475085 and No.12321005. PJ is supported by the European Research Council (ERC) under the European Union’s research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP). MM is supported by a Royal Society Newton International Fellowship (NIF/R1/232539). We acknowledge financial support, supercomputing resources and support from ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing – and hosting entity, funded by European Union – NextGenerationEU. We are grateful to each other’s institutions and the Pauli Center for Theoretical Studies for hospitality.

### v2

@@ Abstract
journal: Physics Letters B
[…]
[first]organization=School of Physics, Shandong University,city=Shandong, postcode=250100, state=Jinan, country=China
[…]
[second]organization=Physik-Institut, Universität Zürich,addressline=Winterthurerstrasse 190, city=Zürich, postcode=CH-8057, country=Switzerland
[…]
[third]organization=Institute for Particle Physics Phenomenology, Department of Physics, Durham University,city=Durham, postcode=DH1 3LE, country=UK
[…]
[fourth]organization=Università degli Studi di Milano-Bicocca & INFN,addressline=Piazza della Scienza 3, city=Milano, postcode=20126, country=Italy
1 Introduction
Experiments at present and future colliders plan to deliver an unprecedented amount of data and significantly increase our sensitivity to new physics if equally precise Standard Model predictions are available. In high-energy collisions, the effects of QCD are dominant and their evaluation at higher perturbative orders in the strong coupling $\alpha_{s}$ requires techniques for the identification and removal of infrared (IR) divergences from the constituent scattering amplitudes, known as infrared subtraction schemes.
@@ 1 Introduction
General approaches have been developed for next-to-leading order (NLO) calculations [32, 68], followed by several strategies at next-to-next-to-leading order (NNLO) [78, 17, 45, 51, 31, 46, 47, 71, 22, 27, 97, 88, 103, 13] and ongoing efforts to include higher-multiplicity final states and to generalize the current formalisms [74, 66, 16, 53, 54]. NNLO-accurate predictions are nowadays available for several Standard Model processes, see [89] for a review. N3LO accuracy has recently been reached for color-singlet production and decay processes for both inclusive [5, 56, 98, 57, 33, 34, 7, 61, 60, 59] and differential [38, 23, 37, 36, 101, 44, 35, 15, 100, 64, 87] observables, relying either on $q_{T}$-slicing [31] or the Projection-to-Born [22] method. Neither of these approaches is applicable to arbitrary processes and observables, especially in the presence of final-state jets, making the development of a general N3LO subtraction scheme highly desirable.
@@
This work represents the first N3LO differential calculation obtained with a subtraction scheme that is not restricted to specific classes of processes and does not rely on the prior knowledge of inclusive quantities. Our calculation is therefore a crucial step in tackling more complicated processes with jets at high perturbative order [26] relevant to the physics goals of the upcoming High-Luminosity phase of the LHC [108]. Moreover, results at this level of precision are relevant for future QCD precision studies [104, 52] at the next generation of electron-positron colliders, such as the FCC-ee [2, 1, 10] and CEPC [4, 42, 3].
2 Method
In the following, we outline the extension of antenna subtraction for final-state radiation to N3LO. The $\mathcal{O}(\alpha_{s}^{3})$ correction to two-jet production at electron-positron colliders receives contributions from final states with up to five partons. We denote the contribution to the N3LO coefficient from a configuration with $n$ final-state partons by $\text{d}\sigma^{(3)}_{n}$. The triple-real contribution $\text{d}\sigma^{(3)}_{5}$ is given by three real emissions on top of the Born-level process, the double-real-virtual contribution $\text{d}\sigma^{(3)}_{4}$ by two real emissions at one loop, the real-double-virtual contribution $\text{d}\sigma^{(3)}_{3}$ by one real emission at two loops and finally the triple-virtual contribution $\text{d}\sigma^{(3)}_{2}$ by the three-loop correction. Examples of Feynman diagrams contributing to different types of N3LO corrections are shown in Figure 1.
[…]
$\text{d}\sigma_{5,\text{finite}}^{(3)}$ $=\text{d}\sigma_{5}^{(3)}$ $-\text{d}\sigma^{(3)}_{5,S_{1}}$ $\hskip-2.0pt-\text{d}\sigma^{(3)}_{5,S_{2}}$ $\hskip-4.0pt-\text{d}\sigma^{(3)}_{5,S_{3}}\,$, (1)
$\text{d}\sigma_{4,\text{finite}}^{(3)}$ $=\text{d}\sigma_{4}^{(3)}$ $-\text{d}\sigma^{(3)}_{4,V_{1}}$ $\hskip-2.0pt-\text{d}\sigma^{(3)}_{4,V_{1}S_{1}}$ $\hskip-4.0pt-\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}\,$, (2)
$\text{d}\sigma_{3,\text{finite}}^{(3)}$ $=\text{d}\sigma_{3}^{(3)}$ $-\text{d}\sigma^{(3)}_{3,V_{2}}$ $\hskip-2.0pt-\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}\,$, (3)
@@ 2 Method
$\text{d}\sigma^{(3)}_{4,V_{1}}$ $=$ $-\hskip-2.0pt\int_{1}\text{d}\sigma^{(3)}_{5,S_{1}}\,$, (5)
$\text{d}\sigma^{(3)}_{3,V_{2}}$ $=$ $-\hskip-2.0pt\int_{1}\text{d}\sigma^{(3)}_{4,V_{1}S_{1}}\hskip-2.0pt-\hskip-2.0pt\int_{2}\text{d}\sigma^{(3)}_{5,S_{2}}\,$, (6)
$\hskip-12.0pt\text{d}\sigma^{(3)}_{2,V_{3}}$ $=$ $-\hskip-2.0pt\int_{1}\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}\hskip-2.0pt-\hskip-2.0pt\int_{2}\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}\hskip-2.0pt-\hskip-2.0pt\int_{3}\text{d}\sigma^{(3)}_{5,S_{3}}\,$, (7)
@@
The subtraction terms are built from antenna functions with the following notation: $X_{n}^{\ell}$ is a $n$-parton $\ell$-loop antenna function intended to subtract the singular behavior of n2n-2 unresolved partons between a pair of hard radiators at $\ell$ loops. $X$ is a placeholder for a particular partonic configuration [78]. The quark-antiquark antenna functions relevant for our process are obtained from color-ordered matrix elements for the decay of a photon to $n$ partons [77, 91], denoted by $M_{n,\gamma^{*}}^{\ell}$. The integrated counterpart of an antenna function, $\mathcal{X}_{n}^{\ell}$, is obtained by analytic integration over the inclusive phase space of the n2n-2 unresolved emissions. All the N3LO final-state antenna functions were recently integrated in [91, 40, 39].
@@
where $M_{n}^{0}$ represents a generic tree-level color-ordered matrix element for $e^{+}e^{-}\to n$ partons and the shorthand symbol $\otimes$ encodes a momentum mapping from a higher-multiplicity to a lower-multiplicity phase space, in order to absorb unresolved emissions. For each term, a sum over external partons and color correlations is understood. The jet function $J_{n_{j}}^{(n_{p})}$ reconstructs $n_{j}$ jets from $n_{p}$ final-state partons. The sums in (13a) and (13b) run over configurations where the two unresolved emissions are respectively color-connected ($cc$) and almost color-connected (accacc[78, 81].
@@
Here the $X_{5}^{0}$ antenna functions address color-connected triple-unresolved emissions. In particular, they contain the singular behavior in triple-soft [29, 50, 28] and quadruple-collinear [49, 48] limits. A $5\to 2$ antenna momentum mapping [94] is used to relate the triple-real and the fully resolved phase space, and we apply phase space sectors [41] to avoid partial fractioning the antenna functions. The remaining terms are needed to remove the single- and double-unresolved limits of the $X_{5}^{0}$, which are already subtracted by $\text{d}\sigma^{(3)}_{5,S_{1}}+\text{d}\sigma^{(3)}_{5,S_{2}}$. The five-particle antenna functions mimic the full matrix element in unresolved limits and factorize onto lower-multiplicity antenna functions as follows:
$X_{5}^{0}$ $\xrightarrow[]{\text{single unres.}}$ $X_{3}^{0}\otimes X_{4}^{0}\,$, (15)
$X_{5}^{0}$ $\xrightarrow[cc]{\text{double unres.}}$ $X_{4}^{0}\otimes X_{3}^{0}\,$, (16)
$X_{5}^{0}$ $\xrightarrow[acc]{\text{double unres.}}$ $(X_{3}^{0}\otimes X_{3}^{0}+\mathcal{S}\otimes X_{3}^{0})\otimes X_{3}^{0}\,.$ (17)
@@
3 Calculation
Our calculation was performed with a version of the NNLOJET Monte Carlo framework [90] extended for an N3LO calculation. For the N3LO antenna functions $X_{n}^{\ell}$, we adapted the implementation of the relevant matrix elements in EERAD3 [83]. The necessary helicity amplitudes for $e^{+}e^{-}\to\text{partons}$ are available in NNLOJET as part of the calculation of three-jet production at NNLO [73]. The five-parton tree-level amplitudes used in our calculation were computed in [86, 11, 62], the four-parton one-loop amplitudes in [84, 25, 12] and the three-parton two-loop amplitudes in [69, 70]. Finally, the three-loop quark form factor was computed in [8, 72, 95].
@@ 3 Calculation
4 Results
All our results are normalized by the respective Born-level cross section $\sigma_{0}$ for $e^{+}e^{-}\to q\bar{q}$. As a consequence, the electroweak coupling drops out and our results are valid for both $\gamma^{*}$ and $Z$ as the intermediate boson. We neglect the singlet contribution which is numerically small, as observed for example in [55, 67]. We set $\sqrt{s}=m_{Z}=91.2$ GeV and $\alpha_{s}(m_{Z})=0.118$. However, for the observables we consider, the dependence on the value of the energy scale only enters through the running of the strong coupling. The renormalization scale is chosen to be equal to the center-of-mass energy, $\mu_{r}=\sqrt{s}$, and theory uncertainties are estimated by varying it within $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$.
@@ 4 Results
$\sigma^{(3)}_{2j}(y_{\text{cut}}=0.1)$ $=\sigma_{0}\left(\frac{\alpha_{s}}{2\pi}\right)^{3}(-453\pm 31)\,$, (19)
$\sigma^{(3)}_{2j}(y_{\text{cut}}=0.01)$ $=\sigma_{0}\left(\frac{\alpha_{s}}{2\pi}\right)^{3}(-2584\pm 40)\,.$ (20)
@@
5 Conclusions
We presented the first direct fully differential calculation of jet production in electron-positron annihilation at third order in QCD. We described the structure of the IR counterterms for the final-state radiation in $e^{+}e^{-}\to jj$ within antenna subtraction, including novel triple-unresolved limits. As an application, we computed the production rate of two jets and the distribution of the energy of the leading jet at N3LO at different values of the jet reconstruction parameter.
@@ 5 Conclusions
Acknowledgements
We are grateful to Thomas Gehrmann, Nigel Glover and Aude Gehrmann-de Ridder for support and encouragement to pursue this work. We thank Elliot Fox for comments on the manuscript. XC is supported by the National Science Foundation of China (NSFC) with grant No.12475085 and No.12321005. PJ is supported by the European Research Council (ERC) under the European Union’s research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP). MM is supported by a Royal Society Newton International Fellowship (NIF/R1/232539). We acknowledge financial support, supercomputing resources and support from ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing – and hosting entity, funded by European Union – NextGenerationEU. We are grateful to each other’s institutions and the Pauli Center for Theoretical Studies for hospitality.